OpenStudy (mtalhahassan2):

Conside the function f(X)=x^3-4x^2+4x. a) Estimate the instantaneous rate of change in f(x) at x= 2. b) What does your answer to part a) tell you about the graph of the function at x=2? c) Sketch a graph of f(x) by first finding the zeros of f(x) to verify your answer to part b).

OpenStudy (mtalhahassan2):

so do we have to use this formula: f(a+h) -f(a) divide h

OpenStudy (anonymous):

Yes

OpenStudy (mtalhahassan2):

oh ok

OpenStudy (triciaal):

the rate of change is the slope also the first derivative find the expression then let x = 2 to get the rate at that moment

OpenStudy (mtalhahassan2):

ok

OpenStudy (mtalhahassan2):

so do we have to use this formula: f(a+h) -f(a) divide h

OpenStudy (mtalhahassan2):

@triciaal

OpenStudy (triciaal):

when x = 0 f(0) = 0 when x = 2 f(2) = 2^3 -4*2^2 + 4*2 = 8-16 + 8 = 0 F(X) = Y change in y/ change in x = 0/2 f(x) = x^3 -4 x^2 + 4 x when f(x) = 0 x(x^2 - 4 x + 4 ) = 0 x(x-2)(x-2) = 0 x = 0, x = 2

OpenStudy (triciaal):

yes?

OpenStudy (mtalhahassan2):

so which be the answer

OpenStudy (mtalhahassan2):

x=0 or x=2

OpenStudy (mtalhahassan2):

@triciaal

OpenStudy (triciaal):

the graph crosses the x-axis at 3 places approximately (-1, 0) (1, 0) and (4, 0) maximum at (0, 4) and there is a minimum when x is between 1 and 4

OpenStudy (triciaal):

when x = 2 y = -4

OpenStudy (mtalhahassan2):

but at the back of the book the answer is about 0

OpenStudy (triciaal):

x-2 (x-1)^2 -1 h = 1 what do you mean? the answer for what ? from above we know the rate of change at x = 2 is 0

OpenStudy (mtalhahassan2):

yep

OpenStudy (triciaal):

here is a sketch of the graph